Circular No. 6665

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

              URL http://cfa-www.harvard.edu/cfa/ps/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

SUPERNOVA 1987A IN THE LARGE MAGELLANIC CLOUD C. S. J. Pun, G. Sonneborn, C. Bowers, T. Gull, S. Heap, R. Kim-

ble, S. Maran, and B. Woodgate, Goddard Space Flight Center, NASA, report for the Space Telescope Imaging Spectrograph (STIS) Instrument Definition Team: “Medium-resolution spectral images (range 481.8–510.4 and 629.5–686.7 nm, resolution ∼ 6000) of SN 1987A and its circumstellar ring system were obtained with the Hubble Space Telescope (+ STIS) on Apr. 26. The spectra were taken through a 52′′ × 2′′ slit aligned with the major axis of the inner ring. Spectral images (spatial FWHM ∼ 0′′.10) of the inner ring were obtained in Hα, Hβ, [O III] at 495.9 and 500.7 nm, [O I] at 630.0 and 636.4 nm, [N II] at 654.8 and 658.3 nm, and [S II] at 671.6 and 673.1 nm. The narrow emission lines (FWHM ∼ 10 km/s) from the ring are not resolved spectrally in our data. Emission features from the supernova debris (FWHM ∼ 3000 km/s) are dispersed and weakly detected only in Hα and Hβ. A blueshifted emission feature, prominent in Hα and less so in Hβ and [O I] 630.0-nm, originates on the inner ring at p.a. 31^{o} ± 8^{o} (i.e., in the ring’s northeast quadrant), believed to be part of the ring tilted toward the earth. The Hα emission from this point has a blue emission wing extending to −250 km/s; there is no redshifted emission. This suggests that the blueshifted emission is excited by interaction between high-velocity gas streaming toward us and material in or close to the inner ring. However, the velocity width of this feature is an order of magnitude narrower than that expected from circumstellar interaction models (e.g., Borkowski et al. 1997, Ap.J. 476, L31). Portions of the outer rings (Burrows et al. 1995, Ap.J. 452, 680) are also detected in Hα and [N II] emission. In [N II] 658.3-nm, the portion of the northern outer ring that crosses the supernova debris disappears at the position of the debris in our spectra, where there is a gap of ∼ 0′′.3 in the outer ring, indicating that this section is located behind the debris. The STIS spectral images of SN 1987A in Hα, [N II], [S II], and [O III] may be viewed at

http://oposite.stsci.edu/pubinfo/PR/97/14.html.”

SUPERNOVA 1997bq IN NGC 3147 Unfiltered CCD magnitudes from S. Moretti and S. Tomaselli, Forlí,

Italy, obtained with a 0.35-m telescope: May 2.9 UT, 15.0; 14.93, 15.6 (GSC comparison stars).

1997 May 20  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1997 CBAT  Daniel W. E. Green